OpenStudy (anonymous):

help ? find the focus and directrix (x-2)^2=-20(y+4) a) (-7,-4) x=3 b)(2,-9) y=1 c)(3,-4)x=-7 d)(2,1) y=-9

OpenStudy (michele_laino):

hint: we can rewrite the equation of your parabola as follows: \[\Large y + 4 = - \frac{1}{{20}}{\left( {x - 2} \right)^2}\] now, if we make this traslation: \[\Large \left\{ \begin{gathered} y + 4 = Y \hfill \\ x - 2 = X \hfill \\ \end{gathered} \right.\] where X, Y is the new reference system, we got: \[\Large Y = - \frac{1}{{20}}{X^2}\]

OpenStudy (anonymous):

so would it be b ?

OpenStudy (michele_laino):

the focus of my parabola is located at: \[\Large X = 0,Y = - 5\]

OpenStudy (michele_laino):

now if I substitute those values, I get: \[\Large \left\{ \begin{gathered} y + 4 = - 5 \hfill \\ x - 2 = 0 \hfill \\ \end{gathered} \right.\] please solve for x, and y

OpenStudy (anonymous):

y=-9 x=2

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

thank you !

OpenStudy (michele_laino):

the directrix is: Y=5 so we have: \[y + 4 = 5\]